OpenStudy (anonymous):

Write in simplified radical form by rationalizing the denominator √10 - √ 3 / √10 +√3 Would this just be 0 ? :-/

OpenStudy (anonymous):

(√10 - √ 3) / (√10 +√3) Like this?

OpenStudy (anonymous):

yup x-actly

OpenStudy (anonymous):

Good. (√10 - √ 3) / (√10 +√3) = (√10 - √ 3) / (√10 +√3) * (√10 - √ 3) / (√10 - √3) =(100-20√3 +9) / (100-3) = (109-20√3)/ 97

OpenStudy (radar):

\[(\sqrt{10}-\sqrt{3}) \times (\sqrt{10}-\sqrt{3})\over(\sqrt{10}+\sqrt{3}(\sqrt{10}-\sqrt{3})\] is equal to \[10-2\sqrt{30}+3\over 10-3\] this then simplifies to\[13-2\sqrt{30}\over 7\]

OpenStudy (radar):

left out a closing paren in the denominator but does not affect the solution